$SPAD/src/input c06ekf.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)clear all
showArrayValues true
showScalarValues true
x:Matrix SF:=
 [[1.00 ,1.00 ,1.00 ,1.00 ,1.00 ,0.00 ,0.00 ,0.00 ,0.00 ]]
y:Matrix SF:=
 [[0.50 ,0.50 ,0.50 ,0.50 ,0.00 ,0.00 ,0.00 ,0.00 ,0.00 ]]
n:=ncols x
job:=1
c06ekf(job,n,x,y,-1)
job:=2
result:=c06ekf(job,n,x,y,-1)

———-

References^{[1] nothing}

4